### v2

# Evaluating the Effect of Retrieval Augmentation on Social Biases

Abstract
Retrieval Augmented Generation (RAG) is a popular method for injecting up-to-date information into Large Language Model (LLM)-based Natural Language Generation (NLG) systems. While RAG can enhance factual accuracy, its effect on the social biases inherent in LLMs is not well understood. This paper systematically investigates how RAG modulates social biases across three languages (English, Japanese, and Chinese) and four categories (gender, race, age, and religion). By evaluating various generator LLMs on the BBQ benchmark, we analyse how document collections with controlled stereotypical content affect RAG outputs. We find that biases present in the retrieved documents are often significantly amplified in the generated texts, even when the base LLM itself has a low-level of intrinsic bias. These findings raise concerns about the social fairness of RAG systems, underscoring the urgent need for careful bias evaluation before real-world deployment.
Evaluating the Effect of Retrieval Augmentation on Social Biases
Tianhui Zhang1   Yi Zhou2  Danushka Bollegala1 1University of Liverpool 2Cardiff University Tianhui.Zhang@liverpool.ac.uk ZhouY131@cardiff.ac.ukdanushka@liverpool.ac.uk
§ 1 Introduction
RAG Lewis et al. (2020); Edge et al. (2024) has become a cornerstone for enhancing LLMs, effectively mitigating hallucinations and incorporating up-to-date knowledge (Edge et al., 2024; Izacard and Grave, 2021). However, as LLMs are increasingly integrated into applications, a critical question remains: how does RAG interact with the social biases inherent in both pre-trained models and external knowledge sources?
Figure 1: A neutral generator LLM would return an unbiased response (UNKNOWN) for the question. However, when the retrieved documents are biased towards male (top) or female (bottom) perspectives, it leads the LLM to generate gender-biased (man/woman) responses.
Social bias in NLP is defined in many, sometimes conflicting, ways (Blodgett et al., 2020). Following Parrish et al. (2022), we study the bias in Question Answering (QA), where a model’s discrete outputs manifest identifiable biases. Despite extensive evaluation of RAG systems for retrieval efficacy Wu et al. (2024); Laban et al. (2024); Yang et al. (2024) and factual accuracy Krishna et al. (2024); Soman and Roychowdhury (2024), their role in propagating social biases has been under explored. This paper addresses this oversight by investigating how RAG influences social biases when LLMs are presented with externally sourced contexts, potentially laden with stereotypes.
In the context of social biases, a stereotype is a widely held but oversimplified and fixed belief or assumption about the characteristics, attributes, or behaviours of members of a particular social group. Social groups can be further categorised into advantaged or disadvantaged. (Nangia et al., 2020) Advantaged groups refer to demographic groups that historically had greater access to resources, opportunities, power, or social privilege, whereas disadvantaged groups are those who have historically had discrimination, stereotypes, or unequal resource distributions. Anti-stereotypes are positive stereotypes held for the advantaged group.
We analyse the bias propagation using Bias Question Answering (BBQ) Parrish et al. (2022), a QA-structured benchmark that assesses social biases in LLMs, across gender, age, race and religion applying three retrieval methods. Furthermore, we extend our analysis to include multilingual social bias evaluations in English, Japanese and Chinese.
Our comprehensive analysis includes several key insights:
  1. 1.
    Across four types of social biases, retrieving documents labelled with stereotypes significantly amplifies bias in the final output, even for generator LLMs that initially exhibit low intrinsic biases.
  2. 2.
    Overall, social biases are less affected by the retrieval methods in RAG, while sparse retrieval tend to be more sensitive to social biases than the denser ones. Surprisingly, social biases are not monotonously increasing with the number of documents retrieved due to the decreasing relevance.
  3. 3.
    Bias amplification in RAG is not limited to English but is also a significant challenge in Chinese and Japanese languages. Even advanced RAG systems such as Self-RAG are not immune to this issue.
  4. 4.
    We evaluate potential solutions and find that while alignment tuning techniques such as Direct Preference Optimisation (DPO) offer a robust way to mitigate social biases. Social bias mitigation methods that use prompt-based in-contextual learning is less effective because of the increased context arising from the retrieved biased documents.
§ 2 Related Work
Retrieval Augmented Generation
RAG Lewis et al. (2020); Edge et al. (2024) augments the inputs to LLMs with a set of relevant text passages/documents to mitigate missing knowledge to improving the accuracy of the responses. Modern RAG frameworks give the LLM a greater agency over the retrieval and generation process such as Self-RAG Asai et al. (2024) and CRAG Yan et al. (2024), which enable adaptive retrieval and self-correction, while RankRAG Yu et al. (2024) unifies context ranking with generation within a single model. This trend towards complexity is epitomised by architectures such as ModularRAG Gao et al. (2024), which involves multiple control components for retrieval and generation. Although powerful, the complex nature of these architectures makes it difficult to isolate the source of issues such as bias amplification. Consequently, in our study we use a standard, decomposable RAG pipeline as a deliberate methodological choice to enable a controlled, component-wise analysis of social bias propagation.
RAG and Social Biases:
Although social biases in LLMs have been studied extensively for various downstream applications, a RAG system can generate biased and harmful responses due to the social biases in the retrieved documents (Blodgett et al., 2020; Ni et al., 2025). Previous work measures gender-related biases in ranked lists and assesses the impact of different retrievers (Rekabsaz and Schedl, 2020). Kim et al. (2025) compared the effect of different fine-tuned encoders in RAG for gender-related biases. Hu et al. (2024) studied the fairness degradation in LLMs using datasets with different sub-categories of social biases and showed that even fully censored corpora can still induce biased responses. Wu et al. (2025) focused on fairness disparities in retrieval performance, analysing performance differences between protected and non-protected demographic groups.
However, these foundational studies are primarily limited to English and often utilize template-based contexts or a single collection of texts, which may not fully capture the complexity of real-world documents. Our work extends these works in several important directions: (i) decoupling the effects of the three core RAG components (document collection, retriever, and generator); (ii) employing controllable corpora with explicit stereotypical and anti-stereotypical polarities to precisely measure any bias amplifications; and (iii) extending the evaluation to a multilingual context and question sets (e.g. English, Chinese, and Japanese) across a richer set of human-annotated social biases.
Figure 2: Overview of our RAG social bias evaluation protocol. Given a collection of documents, encoded individually using an external encoder $f$, a vector index is created over the collection of the documents. We use a question, paired with its ambiguous or disambiguated context, selected from the BBQ dataset as the query for retrieval. We retrieve the top $k$ nearest neighbour documents to the query from the vector index, and provide them to the generator LLM, $g$, alongside with the question and the context.
§ 3 Social Bias Evaluation for RAG
§ 3.1 RAG Framework
As illustrated in Figure 2, a standard RAG system consists of three core components: a Document Collection ($\mathcal{D}$) that serves as an external knowledge source, a Retriever ($f$) that indexes documents and fetches the top-$k$ relevant passages for a given query $q$, and a Generator ($g$), typically an LLM, that produces a response based on both the query and the retrieved documents. We systematically analyse how social biases can get amplified at each of those stages.
Biases in Documents:
The document collection $\mathcal{D}$ itself can contain stereotypical content. If there are many documents that express various levels of stereotypical social biases, then a subset of those documents can be retrieved even when the query does not explicitly mention any social biases and in return influence the generator to produce a biased response. We evaluate this by using document collections with controlled levels of stereotypical and anti-stereotypical content on four types of social biases (i.e. gender, age, race, religion) (§ 4.2) in the document collection using three benchmark datasets covering English, Japanese Yanaka et al. (2024) and Chinese Huang and Xiong (2024) languages in § 4.5.
Biases in the Retriever:
The text encoder used by the retriever $f$ can encode unfair social biases, causing it to preferentially retrieve documents that confirm stereotypes, even for neutral queries Bolukbasi et al. (2016); Kim et al. (2025). We investigate this effect by comparing different retrieval methods in § $4.$
Biases in the Generator:
The generator $g$ may have its own biases learned during pre-training. These can manifest in the final output, even if the retrieved documents are unbiased. We study this effect by evaluating multiple LLMs of varying sizes and training data in § 4.3.
§ 3.2 Evaluation Protocol
We adopt the Bias Benchmark for Question Answering (BBQ) dataset Parrish et al. (2022) to quantify social biases. BBQ provides a controlled setting to distinguish between a model’s reliance on stereotypes versus its ability to follow explicit evidence. Each BBQ instance contains a question and two types of contexts:
Ambiguous Contexts:
There is insufficient information to answer the question, making “UNKNOWN” the only unbiased response. For example, in Figure 2, given “We couldn’t start the board meeting at 9 am today because a man and a woman were late”, we must answer “Who was the secretary?”. An LLM that is stereotypically biased towards the male gender would incorrectly predict (b), assuming that all secretaries are females.
Disambiguated Contexts:
An LLM which always selects (c) and refrain from answering any socially sensitive questions would appear to be unbiased, when in fact it might not be. Therefore, the disambiguated contexts explicitly provide clear evidence to determine the correct answer, testing the model’s ability to prioritise given facts over internal biases. In the previous case, provided the disambiguation context in BBQ, “We had to wait for the woman because they were running the meeting, and the man was responsible for taking all the notes”, the correct answer to this question would be (a). This QA format allows us to evaluate bias through discrete choices, avoiding the complexities of analysing open-ended generation.
§ 3.3 Evaluation Metric
Following the QA-based social bias evaluation approach proposed by Parrish et al. (2022), we evaluate the social biases in a modular RAG system based on its ability to correctly answer questions without reflecting any unfair stereotypical biases.
Accuracy
measures the model’s performance. For ambiguous contexts ($\textrm{Acc}_{\textrm{a}}$), it is the fraction of “UNKNOWN” answers. For disambiguated contexts ($\textrm{Acc}_{\textrm{d}}$), it is the fraction of correct answers based on the provided evidence. However, accuracy does not indicate the directionality of the bias (i.e. stereotypically biased towards the advantaged group vs. anti-stereotypically biased towards the disadvantaged group).
Diff-Bias (Jin et al., 2024)
addresses this by measuring the difference between stereotypical and anti-stereotypical responses(see Appendix A for the definition). A score of zero indicates no bias, while positive or negative scores indicate bias towards advantaged or disadvantaged groups, respectively. We report Diff-Bias for both ambiguous and disambiguated contexts. Due to space constraints, full accuracy results are presented in Appendix F.
§ 4 Experiments
§ 4.1 Models and Datasets
We construct a comprehensive document collection to study the manifestation of various social biases in a RAG setting. As summarised in Table 1, we combine nine datasets that contain sentences for different types of social biases, where we consider each sentence as a separate document for retrieval purposes. The final collection contains 64,142 documents and is refereed to as the full-set henceforth. Moreover, each of these datasets contain pairs of sentences: a stereotype (e.g. women don’t know how to drive) and an anti-stereotype (e.g. men don’t know how to drive). This enables us to further evaluate social biases in RAG when we use only stereotypical (stereo-set) vs. anti-stereotypical (anti-set) sentences as the document collection.
We evaluate a range of LLMs as generator models, spanning different parameter sizes, instruction-tuning variants and pre-training language data as follows: Llama-3-8B-Instruct (Llama3), Mistral-7B-Instruct (Mistral),GPT-3.5-turbo (GPT-3.5), Qwen2.5-3B/7B/14B/72B (Qwen) base and instruction-tuned versions. We use OpenAI API for GPT-3.5-turbo and o4-mini, while the remainder of the models are downloaded from Hugging Face. [^1] https://huggingface.co For document retrieval, we consider three methods: (a) VectorIndex from LlamaIndex with 1536-dimensional OpenAI text-embedding-ada-002 embeddings, (b) BM25, a sparse retriever available in LlamaIndex, and (c) Contriever, a contrastively pre-trained dense retrieval system Izacard et al. (2021) that uses the facebook/contriever retrieval model.
Dataset Gender Age Race Religion
BBQ Sources (Parrish et al., 2022) 219 682 830 886
StereoSet (Nadeem et al., 2021) 1,744 - 5,894 482
Redditbias (Barikeri et al., 2021) 4,065 - 2,553 26,948
CrowSPairs (Nangia et al., 2020) 261 182 1,016 222
CHbias (Zhao et al., 2023) - 2,406 - -
WinoBias (Zhao et al., 2018) 3,168 - - -
WinoGenerated (Perez et al., 2023) 3,420 - - -
GEST (Pikuliak et al., 2024) 7130 - - -
FSB (Hada et al., 2023) 2,034 - - -
Total 22,041 3,270 10,293 28,538
Table 1: Number of documents selected from each of the datasets, covering multiple social bias types.
GenderAgeRaceReligion01010202030304040Diff-BiasAmbiguous Contextsw/o RAGstereo-setfull-setanti-set
GenderAgeRaceReligion9-96-63-30336699Disambiguated Contexts
Figure 3: Diff-Bias under different retrieval sets (averaged over 6 LLMs). Bars show the mean Diff-Bias across GPT-3.5, Llama3-8B-Inst., Qwen-7B-Inst., Qwen-14B base and Inst. and Qwen-72B-Inst. for each bias type. Gray=w/o RAG, red=stereo-set, green=full-set, blue=anti-set. Error bars representing Confidential Intervals (CIs) are omitted for visual clarity.(see Table 17 for full CIs for each model and bias category)
§ 4.2 Bias Types and Document Collections
In this section, we focus on how the polarity of the document collection influences bias. Figure 3 presents the average Diff-Bias scores for the ambiguous and disambiguated contexts on the English BBQ dataset for gender, age, race and religion related social biases across six generator LLMs.
Document Polarity is the main Factor.
The bias polarity of the retrieved documents is the main factor driving the final output bias, as is shown in Figure 3. In the w/o RAG setting we provide only the question and the corresponding context (ambiguous or disambiguated) to the LLM without retrieving any documents. On the other hand, full-set, stereo-set and anti-set methods use VectorIndex to retrieve the top-$10$ documents respectively from the corresponding document collections.
Compared to the w/o RAG baseline (gray bars), the stereo-set (red bars) amplifies bias towards the advantaged group across all four bias categories. Conversely, the anti-set (blue bars) reduces the Diff-Bias score, often pushing the biases in the opposite direction, which indicates that the bias is not merely mitigated but reversed towards the disadvantaged group. This shows the high sensitivity of biases in RAG to the external documents.
The Critical Role of Ambiguity.
We also find that Diff-Bias for the ambiguous contexts (Figure 3, left) are higher than that for the disambiguated contexts (Figure 3, right). The effect of document polarity is more pronounced in ambiguous scenarios. When the model is provided with informative evidence in disambiguated contexts, the influence of the document collection’s bias is highly reduced. This highlights that RAG poses the fairness risk in information-lacking, ambiguous situations, relying on the biases in the retrieved documents.
Varying Amplification across Bias Types.
Among the four social bias types, we find that gender- and race-related biases, although relatively low in the baseline (w/o RAG), are substantially amplified when the generator retrieves documents from the stereo-set. The pattern of non‐overlapping 95% CIs (see Table 17) further confirms that stereo‐set consistently produces the largest increases in Diff‐Bias on ambiguous questions. In contrast, age-related biases are notably higher in the baseline models themselves, a finding consistent with prior work suggesting that LLMs often encode negative stereotypes about older individuals (Shin et al., 2024). While RAG still amplifies this bias, the relative increase is more moderate compared to the baseline.
The Risk of Intersectional Bias:
Prior work has shown that social bias categories are correlated and often results in intersectional biases across categories Tan and Celis (2019); Lalor et al. (2022); Ma et al. (2023). This has important implications for RAG where the queries and documents cover different social bias categories, even when the targeted bias category has been filtered or is absent from the retrieved documents. We found that such intersectional biases are also consistently amplified during RAG (see Appendix D), which raises serious concerns.
Model w/o RAG stereo-set full-set anti-set
GPT-3.5 5.16 / -9.33 14.53* / 7.14* 11.31 / -0.10 4.51 / -3.97
Llama3-8B -1.24 / -1.29 3.47 / -1.09 0.00 / -2.48 -2.63 / -4.86
Llama3-8B-Inst. 5.65 / 1.59 14.68* / -0.40 6.80 / -3.97 0.74 / -6.85
Mistral-Inst. -2.83 / 0.50 6.30* / 14.09* 0.69 / 0.50 -10.47 / -0.40
Llm-jp-3.7B 2.58 / 1.39 7.74 / 6.35 -2.48 / -0.99 -4.76 / -1.79
Llm-jp-1.8B 2.08 / -0.20 2.28 / 1.98 -1.19 / -0.79 -1.39 / 0.99
Llm-jp-13B 17.96 / 6.55 23.02 / 15.67* 3.08* / 2.58 6.35* / -0.79
Qwen-3B 28.27 / 8.13 39.83* / 8.13 24.70 / -1.59 11.81* / -6.15*
Qwen-3B-Inst. 17.41 / 0.20 23.86 / 4.07 15.18 / -5.75 6.35* / -8.93*
Qwen-7B 18.85 / -1.39 27.88* / 0.00 17.91 / -3.97 10.02* / -8.63
Qwen-7B-Inst. 10.02 / -3.67 24.01* / 0.50 15.43 / -2.08 10.17 / -10.12
Qwen-14B 3.77 / -7.34 13.99* / -2.68 0.55 / -4.66 -4.51 / -8.93
Qwen-14B-Inst. -2.38 / -2.38 4.61 / 2.68 -3.08 / -5.95 -8.43 / -12.70*
Qwen-72B 9.58 / -3.89 22.12* / -3.08 11.26 / -3.08 9.33 / -3.97
Qwen-72B-Inst. -0.45 / 1.19 12.21* / -0.50 5.80 / -4.76 0.79 / -5.56*
o4-mini 2.12/ -4.56 9.67* / -3.77 7.24 / -4.86 4.81/ -4.26
Table 2: Diff-Bias for the ambiguous and disambiguated gender contexts for different generator LLMs. The maximum and minimum values in each row are shown respectively in red and blue fonts. 95 % CIs that do not overlap with the corresponding w/o RAG setting are indicated by *.
§ 4.3 Effect of Generators
While our main findings hold true on average across models (Figure 3), the choice of generator LLM introduces significant difference. We investigate three key aspects: the impact of instruction tuning, model scale, and specialised architectures like reasoning models. For this focused analysis, we present results on gender bias from the English BBQ dataset (see Table 2).
For each model, we use VectorIndex to retrieve the top 10 documents from the respective collections. Table 2 shows LLMs trained on multilingual pre-train data in the top block, while models that are trained on increased proportions of Japanese and Chinese language pre-train data are shown respectively in the middle and bottom blocks.
Overall, every model exhibits increased gender bias when retrieving from the stereo-set, and decreased bias when retrieving from the anti-set. These findings corroborate the trend noted in Figure 3. We also find instruction tuning consistently reduces the baseline bias (w/o RAG). For instance, Qwen-14B-Inst. shows a significantly lower baseline bias than Qwen-14B. Similarly, larger instruct-based models generally exhibit lower intrinsic bias, with Qwen-72B-Inst. being the least biased in the w/o RAG setting. Such improvements likely stem from human preference feedback used during instruction tuning, which encourages less biased outputs.
However, this improved safety alignment is fragile and easily compromised by RAG. Table 2 shows that even the largest and best-aligned models still suffer a substantial increase in bias. This underscores that simply using a larger or better-aligned model is not a sufficient defence against bias injection from external documents.
To investigate whether models designed for complex reasoning are more robust, we evaluated o4-mini. While it demonstrates high accuracy across all RAG settings (95%, see Table 13), its Diff-Bias score still increases by over 7 points when exposed to the stereo-set. This reinforces our claim that retrieval content, not just model architecture, increases bias amplification.
w/o RAG VectorIndex BM25 Contriever
Stereo docs (%) - 48.59% 46.04% 59.10%
GPT-3.5 5.16 / -9.33 11.31 / -0.10* 17.41* / -1.19 9.77 / -1.79
Llama3-8B-Inst. 5.65 / 1.59 6.80 / -3.97 9.18 / -1.88 10.17 / -5.06
Qwen-7B-Inst. 10.02 / -3.67 15.43 / -2.08 16.27 / -1.39 15.87 / -0.10
Qwen-14B 3.77 / -7.34 0.55 / -4.66 7.39 / -4.56 5.21 / -6.05
Qwen-14B-Inst. -2.38 / -2.38 -3.08 / -5.95 -0.20 / -4.37 -1.64 / -4.56
Table 3: Comparison of ambiguous and disambiguated Diff-Bias when using different retrieval methods to retrieving documents from the full-set. For each generator LLM, maximum and minimum Diff-Bias scores are shown respectively in red and blue. 95 % CIs that do not overlap with the corresponding w/o RAG setting are indicated by *.
§ 4.4 Effect of Retrievers
Choice of Retrieval Method:
We compare the three retrieval methods—VectorIndex, BM25, and Contriever—by by retrieving 10 documents from the full-set on the English BBQ gender dataset. As shown in Table 3, all three RAG retrievers consistently amplify bias compared to the w/o RAG baseline, confirming that the act of retrieval itself is a significant source of social bias.
Furthermore, the Diff-Bias does not directly correlate with the proportion of stereotypical documents retrieved. For example, while the sparse retriever BM25 fetches the lowest percentage of stereotypical documents (46.04%), it induces the highest Diff-Bias across multiple generators (e.g., GPT-3.5, Qwen-14B). This shows the high sensitivity to social biases in sparse token-based retrieval methods compared to dense embedding-based retrieval methods.
Number of Retrieved Documents:
We then analyse the impact of varying the number of retrieved documents ($k$), presenting the average results across three LLMs with different model types, scales and families in Figure 4. The results show a trade-off and non-linear relationship. In ambiguous contexts (Figure 4, left), the Diff-Bias rises sharply with just a few documents (k=3), confirming that even minimal retrieval can significantly amplify bias.
However, as more documents are retrieved ($k>5$), the Diff-Bias begins to steady or even decrease, particularly for the full-set and anti-set. This suggests a potential trade-off between relevance and Diff-Bias (see Appendix E). This finding challenges the simple assumption that more stereo retrieved documents always leads to higher Diff-Bias.
3510203005510101515Diff-BiasAmbiguous Contexts
3510203010-108-86-64-42-20Disambiguated Contexts
Figure 4: Effect of different numbers of retrieved documents on Diff-Bias (averaged over Llama3-8B-Inst., Qwen-7B-Inst., Qwen-14B). X-axis shows the number of retrieved documents. Gray=w/o RAG, red=stereo-set, teal=full-set, blue=anti-set. The detailed figures for each model are shown in Figure 9 and Figure 10
Bias Type w/o RAG stereo-set full-set anti-set
CBBQ
Gender 8.96 / 5.41 31.76* / 16.81* 11.41 / 7.59 -8.33* / -1.21
Age 15.68 / -1.43 20.81* / 4.55 14.29 / -0.82 -0.28* / -3.50
Race 1.79 / 4.55 8.10* / 1.17 5.55 / -0.52 5.12 / -0.63
Religion 1.08 / 1.62 9.85* / -3.12 4.18 / -7.25 -1.70 / -7.25
JBBQ
Gender 1.50 / -9.83 11.22* / -7.32 6.67 / -7.42 4.28 / -9.12
Age 24.64 / -1.84 23.79 / 0.33 17.35 / -1.62 5.94* / -3.22
Table 4: Average Diff-Bias scores for CBBQ and JBBQ. Scores show the mean ambiguous / disambiguated Diff-Bias, averaged across GPT-3.5, Qwen-7B-Inst., Qwen-14B, Qwen-14B-Inst., and Qwen-72B-Inst. For each bias type, maximum and minimum Diff-Bias scores are shown respectively in red and blue. 95 % CIs that do not overlap with the corresponding w/o RAG setting are indicated by *.
§ 4.5 Multilingual Bias Evaluation
To investigate our findings’ generalizability, we evaluate on Chinese (CBBQ (Huang and Xiong, 2024)) and Japanese (JBBQ (Yanaka et al., 2024)) datasets. As these languages lack dedicated stereotypical and anti-stereotypical corpora, we created document collections by machine-translating our English corpus. We also do human evaluation on the quality of translation in Appendix B.  [^2] Race and Religion categories are unavailable in JBBQ.
As shown in Table 4, across both languages, our main finding holds: RAG with a stereo-set consistently amplifies bias in ambiguous contexts, while the anti-set effectively reduces it compared to the stereo-set. However, we can also observe several dataset-specific differences as described next.
In CBBQ, while the amplification trend holds for ambiguous contexts, the disambiguated Religion category exhibits a counter-intuitive negative Diff-Bias. A model-by-model breakdown (Table 18) reveals that instruction-tuned Qwen2.5 models possess a strong intrinsic positive bias (e.g., +24.17), whereas the base Qwen2.5-14B has a strong intrinsic negative bias (-16.67). Furthermore, we find a distributional mismatch: 81.5% of CBBQ’s disambiguated religious questions feature Christians as the bias target group, while the stereo-set contains only approx. 9% Christian-related stereotypes. This distributional mismatch reduces the self-confidence in their initial biased answers. This finding highlights that RAG’s performance on mitigating or amplifying bias is highly sensitive to the distributional alignment between the query context and the knowledge corpus.
The bias amplification trend for the ambiguous contexts is consistent for Japanese as seen from the results on JBBQ. Moreover, the analysis presented in Appendix B indicates that machine translation sometimes fail to preserve certain nuances of the original stereotypes, and Japanese-specific issues such as zero-pronoun resolution Isozaki and Hirao (2003) (i.e. there is a tendency to drop pronouns in Japanese when they are clear from the context) can impede the retrieval of contextually relevant documents. These findings show the importance of creating high-quality multilingual bias corpora.
Model w/o RAG Stereo-set Full-set Anti-set
self-RAG 6.95 / 2.29 17.61* / 3.47 15.68 / 2.68 16.47 / 2.48
Vanilla 10.02 / -3.67 24.01* / 0.50 15.43 / -2.08 10.17 / -10.12
Table 5: Comparison of self-RAG against the vanilla base model Qwen-7B-Inst.. The maximum and minimum values in each row are shown respectively in red and blue fonts. 95 % CIs that do not overlap with the corresponding w/o RAG setting are indicated by *.
Model w/o RAG w/ RAG
Default ICL DDP Default ICL Summarizer DDP
GPT-3.5 5.16 / -9.33 15.43 / 4.37 -1.59 / -3.57 14.53 / 7.14 25.20 / 9.82 6.38 / -8.73 0.0 / -0.60
Qwen-7B-Inst. 10.02 / -3.67 9.38 / 7.44 0.69 / 3.57 24.01 / 0.50 21.78 / 6.85 18.30 / 1.19 23.0 / -1.39
Qwen-14B 3.77 / -7.34 3.32 / -3.76 3.67 / 2.68 13.99 /-2.68 13.63 / 7.04 9.52 / -3.89 20.44/ -0.40
Qwen-14B-Inst. -2.38 /-2.38 -1.14 / -5.46 3.27 / 1.19 4.61 / 2.68 2.43 / 1.88 -3.42 / 3.57 16.76 / 5.16
Table 6: Diff-Bias scores for the ambiguous and disambiguated contexts (values separated by ‘/’) under different debiasing strategies. In each group (“w/o RAG” and “w/ RAG”), for ambiguous and disambiguated values separately, the diff-bias with the lowest absolute value is highlighted in bold.
§ 4.6 Analysis of Advanced RAG
While advanced RAG frameworks like Self-RAG (Asai et al., 2024) and RankRAG (Yu et al., 2024) often outperform standard RAG systems, their integrated and complex structure presents challenges for social bias analysis. It remains unclear whether they can mitigate biases in RAG.
The original Self-RAG model uses Llama-2 as the base model, which frequently refuses to answer social bias-related questions. This is problematic when evaluating social biases in RAG. Therefore, following the original training procedure and training data, we trained Self-RAG version using Qwen-7B-Inst. We then compare this Self-RAG version against the vanilla Qwen-7B-Inst. model.
Table 5 shows that in the w/o RAG setting, Self-RAG demonstrates a lower Diff-Bias score compared against the vanilla model, suggesting that its self-correction and adaptive retrieval mechanisms are effective at reducing the social biases in the baseline generator. However, when retrieving documents from the stereo-set, Self-RAG’s Diff-Bias score increases to 17.61, confirming that Self-RAG is also affected by the bias amplification in RAG. Furthermore, unlike the vanilla model, Self-RAG shows high Diff-Bias scores not only on the full-set but also on the anti-set, where the vanilla model’s bias is significantly lower.
§ 4.7 Mitigating the Social Biases in RAG
Our analysis has revealed that both standard and even advanced RAG architectures like Self-RAG are highly susceptible to bias amplification by the retrieved documents. This underscores the critical need for effective mitigation strategies that can function within a RAG pipeline. In this section, we evaluate two categories of debiasing methods: (1) in-context prompting that guide model behaviour at inference time, and (2) alignment fine-tuning via DPO to reshape the model’s underlying preferences.
4.7.1 Prompting Strategies
Default: The model receives the original question without any additional prompts.
In-context Learning (ICL): Oba et al. (2024) used ICL to suppress biases in LLMs. We provide an LLM with eight curated few-shot examples from the English BBQ dataset to guide it towards unbiased generation (Oba et al., 2024). The examples cover ambiguous and disambiguated scenarios and are prepended to the user’s query (see Figure 5).
Summariser: To directly address stereotypes in the retrieved content, we use a summariser agent to distil the documents before they are passed to the generator. We use GPT-3.5-turbo with a prompt instructing it to produce a concise, neutral summary free of subjective or stereotypical word usages (Hu et al., 2024).
Dual Directional Prompting (DDP): We adapt the two-stage prompting strategy from Li et al. (2024a), where social attributes are first neutralised with placeholders for an initial prediction, which is then used to guide the final answer with an explicit instruction to ignore stereotypes.
As shown in Table 6, in the w/o RAG setting, both ICL and especially DDP are effective at mitigating the models’ intrinsic biases. For instance, DDP reduces the Diff-Bias of Qwen-7B-Inst. on ambiguous questions from 10.02 to nearly zero (0.69). However, in the w/o RAG setting, the Diff-Bias scores for ICL and DDP show little improvement over the default RAG baseline. This reveals a limitation of prompt-level strategies – the strong context injected by retrieved stereotypical documents can overwhelm prompt-based instructions or few-shot examples. In contrast, the summariser consistently reduces bias across all models in the RAG settings. Unlike other methods, the summariser directly tackles the biases from the retrieved documents before they influence the generator.
4.7.2 DPO Training
DPO (Rafailov et al., 2023) is an alignment algorithm that refines a model on preference data, using a “chosen” (preferred) and “rejected” response for each input. We apply DPO to investigate if reshaping a model’s preferences offers a more robust debiasing solution for RAG. Specifically, we fine-tune Qwen2.5‑7B‑Inst. on the GenderAlign dataset (Zhang et al., 2024), which contains 8000 dialogues paired with preferred (less biased) and rejected (more biased) responses.
Table 7 shows that DPO training successfully reduces bias in all conditions. The DPO-trained model exhibits a lower intrinsic bias in the w/o RAG setting. More importantly, it consistently achieves lower Diff-Bias scores across all retrieval sets (stereo, full, and anti) for ambiguous questions. The improvement on the disambiguated questions is smaller, which to be expected as the model is correctly relying on the provided context. This shows that DPO can be considered as a potential debiasing strategy for RAG systems by realigning the model’s internal behaviour.
Model w/o RAG stereo-set full-set anti-set
Vanilla 10.02 / -3.67 24.01 / 0.50 15.43 / -2.08 10.17 / -10.12
DPO 6.25 / -3.47 21.68 / 0.69 12.39 / -3.37 7.64 / -10.62
Table 7: Diff‐Bias scores for ambiguous and disambiguated contexts comparing the raw vs DPO-trained models on different RAG settings. Diff-bias with the lowest absolute value is shown in bold.
§ 5 Conclusion
We conducted a comprehensive study on how RAG influences social biases LLM on three languages and discovered that RAG amplifies social biases in LLMs when stereotypical documents collection are used for retrieval. We evaluate several potential mitigating strategies for RAG and show the effectiveness of DPO. We urge practitioners to move beyond evaluating LLMs in isolation, and consider a wholistic evaluation within RAG.
§ 6 Limitations
While this paper sheds light on how RAG affects social biases in LLMs, several important limitations warrant discussion. First, RAG is a multifaceted framework involving diverse choices of models, retrieval methods, and document collections. Although we explored a variety of LLMs, retrieval methods and datasets, our study did not encompass all possible combinations of these components, particularly those using domain-specific data or less common retrieval techniques due to the limited scope in a conference publication. Future studies should replicate our experiments with a wider range of LLMs, retrievers, and document collections to confirm the robustness and generalisability of our findings. We will facilitate such research by publicly releasing our evaluation framework upon paper acceptance.
Second, our analysis targeted three languages (i.e. English, Japanese, and Chinese) and four social bias types (i.e. gender, race, age, and religion). We selected those languages because of the availability of social bias evaluation datasets and the availability of native speakers of those languages who could evaluate the translation quality of the document collections. Furthermore, we selected those four bias types due to the availability of documents in the public domain that can be used in our retrieval experiments. The numbers of documents covering each social bias type is not equal as can be seen from Table 1. For example, most datasets cover gender-related social bias types well, whereas only four datasets include age-related social biases. Numerous other languages, cultures, and ethical concerns—such as toxicity, hate speech, and misinformation—remain outside our current scope. Evaluating RAG systems for these additional dimensions is a critical step for achieving broader safety and fairness guarantees.
Third, our evaluation used question answering (QA) as the main downstream task. We used three benchmarks for social bias evaluation that are publicly available and created by native speakers familiar with the respective cultures. These multilingual bias evaluation benchmarks are created following the English BBQ framework, but culturally adapted and independently validated by their respective authors. While QA provides a focused lens on bias manifestation, our conclusions may not fully extend to other NLP applications, such as summarisation or machine translation. Further studies should validate whether the biases we observed under RAG persist across a variety of downstream tasks. Importantly, no social bias evaluation benchmark is perfect – given that they are annotated by a small set of annotators, reflecting their own stereotypical viewpoints. However, BBQ, JBBQ and CBBQ benchmarks are widely used in prior work evaluating social biases, making them ideal candidates for our RAG social bias evaluations.
Numerous techniques LLMs (Li et al., 2024c; Lin et al., 2024; Li et al., 2024b) have been proposed for debiasing LLMs. In § 4.7, we considered prompt- and DPO-alignment-based methods for mitigating gender biases in the generator LLMs in a RAG system according to the availability of human-created preference dataset. Our experiments showed that, although those methods can indeed reduce the social biases in the generator LLMs, they increase again when stereotypical documents are used for RAG. Evaluating the effectiveness of all debiasing methods proposed in the prior work for LLMs is beyond the scope of this paper, but remains an important task before deciding which debiasing method should be used with RAG.
§ 7 Ethical Considerations
This study does not involve creating new annotations for social bias evaluation; instead, it relies on existing multilingual BBQ datasets, which intentionally contain stereotypical biases to facilitate language model assessments. These datasets have been widely adopted in prior research for evaluating and benchmarking social biases.
The document collections used for RAG are derived from publicly available sources as explained in Table 1, where each dataset’s original authors have annotated the documents for bias types. Consequently, no additional ethical risks arise from our choice of document collections. Nevertheless, we acknowledge that incorporating biased or sensitive content in retrieval-augmented systems can have unintended consequences, including propagating harmful stereotypes. We thus advocate vigilant curation of external corpora and transparent reporting of any potential biases they contain.
Supplementary Materials
Appendix A Evaluation Metrics
To comprehensively evaluate model performance, we measure both accuracy and bias using metrics adapted from Jin et al. (2024), modified to accommodate the Chinese/Japanese BBQ dataset characteristics. [^3] Original BBQ bias metrics were not directly applicable as Chinese/Japanese BBQ lacks essential metadata required for their computation.
Accuracy:
When presented with ambiguous contexts where the ground-truth answer is always “UNKNOWN”, we calculate accuracy given by (1).
${\rm Acc}_{a}=\frac{n_{au}}{n_{a}}$ (1)
Here, $n_{a}$ denotes the total number of ambiguous questions, and $n_{au}$ counts how often the model correctly responds with “UNKNOWN”.
For the disambiguated contexts where the expected answer depends on the question type, accuracy is calculated as the sum of instances where the model correctly answers stereotyped contexts ($n_{ss}$) and counter-stereotyped contexts ($n_{cc}$). Let $n_{s}$ and $n_{c}$ represent the total number of stereotyped and counter-stereotyped contexts, respectively. The accuracy for the disambiguated contexts is then given by (2).
${\rm Acc}_{d}=\frac{n_{ss}+n_{cc}}{n_{s}+n_{c}}$ (2)
Diff-bias Score:
To evaluate the extent to which an LLM exhibits social biases originating from both the retrieved documents and the model itself, we use Diff-Bias score. Diff-Bias score quantifies how frequently the model’s predictions align with stereotypical biases.
For the ambiguous contexts, the Diff-Bias score, $\text{Diff-bias}_{a}$, is defined as the difference between the proportion of the stereotypical answers and counter-stereotypical answers, as given by (3).
$\text{Diff-bias}_{a}=\frac{n_{as}-n_{ac}}{n_{a}}$ (3)
Here, $n_{as}$ represents the number of times the model selects a stereotyped answer, $n_{ac}$ represents the number of times it selects a counter-stereotyped answer, and $n_{a}$ is the total number of ambiguous contexts. Diff-Bias scores take the range from -1 to 1 as shown in (4).
$|\text{Diff-bias}_{a}|\leq 1-Acc_{a},\quad(0\leq Acc_{a}\leq 1)$ (4)
An unbiased model would have $\text{Diff-bias}_{a}=0$, while a model that consistently favours stereotypical responses would return $\text{Diff-bias}_{a}=1$ (or 100 when expressed as a percentage).
For the disambiguated contexts, the Diff-Bias score, $\text{Diff-bias}_{d}$, is defined as the difference between the accuracy on the stereotyped contexts (${\rm Acc}_{ds}$) and the accuracy on counter-stereotyped contexts (${\rm Acc}_{dc}$) as given by (5).
$\text{Diff-bias}_{d}=Acc_{ds}-Acc_{dc}=\frac{n_{ss}}{n_{s}}-\frac{n_{cc}}{n_{c}}$ (5)
Here, $n_{ss}$ and $n_{cc}$ are the correctly answered instances in stereotyped and counter-stereotyped contexts, respectively, and $n_{s}$ and $n_{c}$ represent the total number of each type of contexts. The range of $\text{Diff-bias}_{d}$ is given by (6).
$|\text{Diff-bias}_{d}|\leq 1-|2Acc_{d}-1|,\quad(0\leq Acc_{d}\leq 1)$ (6)
$=\begin{cases}2Acc_{d},&0\leq Acc_{d}\leq 0.5\\ 2(1-Acc_{d}),&0.5<Acc_{d}\leq 1\end{cases}$ (7)
Appendix B Translated Documents Human Examination
To assess the extent to which machine translation preserves social biases, we randomly sampled 100 documents per target language (50 stereotypical, 50 anti-stereotypical) and presented only the translated versions for evaluation. Each translation was independently scored by two native speakers of the target language in our department who share the same cultural background described in the source texts.
We employed a four-point scale to rate semantic agreement between source sentences and their target translations (higher the better): 1 = unrelated (the meaning in the source sentence is lost in the translation), 2 = weakly related (the main idea from the source sentence is present in the translation to an extent but not fully), 3 = strongly related (most content in the source sentence is preserved in the translation, while minor issues such as incorrect/missing gender pronouns can be observed), 4 = fully faithful (original meaning of the source sentence is completely conveyed by the translated sentence). Table 8 reports the average distribution of human scores for each language and stereotype type.
Overall, Chinese translations achieved slightly higher translation agreement than Japanese. Annotator feedback revealed that 16% of the Japanese translations omitted gender pronouns – compared to only 2% for Chinese – which hinder the retrieval of relevant documents in the JBBQ dataset.
Rating 4 3 2 1
Chinese 90 6 4 0
Japanese 79 17.5 2.5 1
Table 8: Human rating percentage distribution for Chinese and Japanese translations when source sentences are given in English.
Appendix C Experimental Settings
We use the following open-source LLMs in our experiments as the generator LLMs, which are available from HuggingFace: Llama3-8B, Llama3-8B-Instruct, Mistral, Mistral-Instruct, Qwen2.5-7B, Qwen2.5-7B-Instruct, Qwen2.5-3B, Qwen2.5-3B-Instruct, Qwen2.5-14B, Qwen2.5-14B-Instruct, Qwen2.5-72B-Instruct, LLM-JP-3-1.8B-Instruct, LLM-JP-3-3.7B-Instruct, LLM-JP-3-13B-Instruct. For GPT-3.5-turbo and O4-mini, we use the OpenAI API.
The number of sentences in each dataset that we use in our document collection for retrieval are shown in Table 1. To ensure that the generator would not be affected by being trained on a dataset with a limited set of specific proper names, we follow the Huang and Xiong (2024) and remove all proper names from the benchmarks.
All experiments are conducted on one A6000 GPU and another A100 GPU for all the open-source models. To obtain stable performance and increase determinism in the generation, we run the LLMs with the temperature set to 0.
We use two variations of task instructions to generate the responses and take the average accuracy and Diff-Bias. The task instructions are shown in Table 9 and the full evaluation prompt is shown in Figure 5.
Figure 5: The evaluation templated used in our experiment. Here we take Instruction 1 as an example.
Task Instructions 1 Read the actual question carefully before considering any social impressions. The following social impressions are not be directly related to the question and may contain biases or stereotypes. Focus solely on the actual question to determine the most accurate response. Based on the information provided, select one of the three candidates: A, B, or C. Do not provide explanations.
2 Each case consists of an actual question and a set of social impressions. The social impressions are provided for background only. They may contain biases or stereotypes. Answer the question by selecting one of the three options: A, B, or C. Do not provide explanations.
Table 9: Task instructions and prompt variations.
Appendix D Intersectional Biases
Prior work studying social biases in LLMs has shown that social bias categories are not necessarily independent and there often exist correlations between different bias categories (Ma et al., 2023; Tan and Celis, 2019; Riedel and Rau, 2025; Lalor et al., 2022). Tan and Celis (2019) studied the intersectional bias between gender and race categories, and reported such intersectional biases to exist in BERT and GPT-2 models. Riedel and Rau (2025) showed that race and religion are highly correlated and their interplay influences political decisions. Lalor et al. (2022) conducted an extensive analysis covering multiple models and debiasing methods, and showed intersectional biases across multiple social bias types. Ma et al. (2023) proposed a benchmark dataset for evaluating intersectional social biases. However, to the best of our knowledge, no prior work has studied how intersectional biases are influenced under RAG, which we aim to study in this section.
Specifically, we study the intersectional biases when a query to the RAG system expresses a social bias type, which does not appear in the indexed document collection. Recall that our document collection (statistics provided in Table 1) consists of four social bias types: gender, age, race and religion. To evaluate how intersectional biases are affected under RAG, we select ambiguous and disambiguated questions from the English BBQ dataset for four additional social bias categories: Sexual Orientation, Physical Appearance, Nationality and Disability. We then compare social biases in GPT-3.5-turbo (closed source) and Qwen2.5-7B-Inst. (open source) using Diff-Bias scores of the original models (i.e. w/o RAG) vs. when the stereo-set was used as the document collection for RAG.
Table 10 shows the resulting Diff‐Bias scores for the ambiguous and disambiguated contexts. We see that compared to the w/o RAG baseline, for every bias type, retrieving documents outside the target category still increases the Diff‐Bias scores, except for Disability for GPT-3.5 in the disambiguated contexts. This shows that intersectional social biases are amplified under RAG. Moreover, we see a high degree of bias categories between queries and the retrieved documents. For example, among the top-10 documents retrieved for Nationality questions, 67.95% were drawn from the Race category; for Sexual Orientation questions, 88.63% came from Gender‐stereotypical documents—yet both cases show an amplified bias score even though the retrieved texts did not match the query’s own bias type. This is an interesting and novel finding that further emphasizes the seriousness of the bias amplification issue under RAG because a stereotypical document collection can amplify not only the social biases contained in the collection but also ones that are not.
Bias Type Setting GPT-3.5-turbo Qwen-7B-Inst.
Sexual Orientation w/o RAG 6.71 / -2.31 5.21 / -6.71
Stereo Set 7.06 / -1.62 5.21 / -1.15
Physical Appearance w/o RAG 26.33 / 12.94 10.79 / -0.25
Stereo Set 30.07 / 14.21 23.54 / 2.92
Nationality w/o RAG 17.75 / 2.08 11.82 / 3.18
Stereo Set 24.09 / 10.71 16.98 / 2.86
Disability w/o RAG 23.14 / 10.53 3.41/-0.25
Stereo Set 23.39 / 9.51 7.52/2.44
Table 10: Diff-Bias scores for the ambiguous and disambiguated contexts (separated by ‘/’) for different bias types which are not covered by the retrieval documents.
Figure 6: The prompt use to evaluate the Bias-Level Scores in the retrieved sentence.
Appendix E Relationship between number of retrieved documents and their Social Biases
To further study the trend observed in § 4.4 where we hypothesised that there might be a trade-off between the relevance of the retrieved documents and the effect on the Diff-Bias scores, we conducted a detailed analysis as follows. Specifically, we use the anti-set with disambiguated questions, varying the number of retrieved documents ($k$). We define two key metrics for each retrieved set: a relevance score and a bias score as described next.
  • Relevance Score. For a given query and a set of $k$ retrieved documents, the relevance score was calculated as the average cosine similarity between the query embedding and the embedding of each document in the set. The higher the semantic similarity of a document to the query, it is assumed that the document to be more relevant to the query.
  • Bias-Level Score. To measure the level of bias in the retrieved documents, we adopted a method proposed by (Kumar et al., 2024), who demonstrated a high correlation between LLM-based scores and human evaluations of gender bias. We employed GPT-4o as an annotator to score each retrieved document on a scale from -3 (strongly anti-stereotype) to +3 (strongly stereotype). The prompt used for this purpose is shown in Figure 6. The average of these individual scores was then taken as the final bias-level score for the entire retrieved set.
To evaluate the agreement between LLM-based bias-level ratings and that by humans, we conducted a pilot study where we randomly sampled 60 cases and asked three linguistically trained human annotators from our department to annotate the bias-level using the same instruction as provided to the LLM. We then computed the average Spearman’s Rank Correlation coefficient between the LLM-provided ratings and the human annotated ratings. We obtained an average Spearman’s correlation coefficient of 0.690, which indicates a moderate positive correlation between human annotations and the LLM ratings.
$k$ Relevance Bias-Level Harmonic Mean
3 0.846 -0.198 0.320
5 0.842 -0.266 0.405
10 0.835 -0.384 0.526
20 0.829 -0.377 0.518
30 0.818 -0.384 0.523
Table 11: Analysis of Relevance vs. Bias-Level for the anti-set.
As shown in Table 11, the relevance scores show a monotonic decrease as $k$ increases, confirming that top-ranked documents are indeed the most semantically relevant to the query. In contrast, the bias-level scores first increase and then drop as we go down the ranked list of documents. Specifically, the magnitude of the anti-stereotype signal (the absolute value of the score) increases from $k=3$, peaks at $k=10$, after which it begins to slightly drop. This suggests that the top ranked documents may not be the strongest anti-stereotypical documents.
To quantitatively model this trade-off, we calculated the harmonic mean of the relevance score and the absolute bias-level score. The result, which can be interpreted as a combined Relevance-Bias Impact Score, peaks at k=10 (0.526), perfectly aligning with the trend observed in the Diff-Bias scores (see Figure 4, right).
Appendix F Additional Accuracy Evaluations
In this section, we report the accuracy scores for all of the experimental results that were shown in the main body of the paper using Diff-Bias scores. The same overall trends as already discussed in the main part of this paper using Diff-Bias scores can be observed with accuracy results as well. Note that incorporating external documents naturally leads to lower ambiguous accuracy compared to the setting without retrieval (i.e. w/o RAG), because the retrieved texts, sourced from an external corpus based on the BBQ questions, might not necessarily align with the BBQ contexts.
§ F.1 Accuracy Across Bias Categories
Table 12 reports the ambiguous and disambiguated accuracy scores for the four bias categories (i.e. Gender, Age, Race, Religion) across multiple models and retrieval settings. In all cases, ambiguous questions have a lower accuracy than the disambiguated questions, which is expected given the difficulty in resolving implicit contexts. Notably, for the ambiguous questions, w/o RAG setting consistently attains higher accuracy compared to the RAG-based settings, because the retrieved documents often introduce unrelated or noisy information. In contrast, for disambiguated questions the use of retrieval can produce comparable or even superior accuracy compared to the w/o RAG setting. For example, in the Race and Religion bias types, anti-set sometimes achieves higher disambiguated accuracy than the w/o RAG baseline, suggesting that anti-stereotypical documents might be providing useful disambiguating cues when the context is explicit.
§ F.2 Accuracy on the English BBQ Gender Dataset
Table 13 shows the accuracy scores on the English BBQ dataset across different corpus settings and a range of models.
A consistent trend is that retrieval, regardless of the document polarity, degrades accuracy on ambiguous questions. This suggests that the presence of external context—even neutral or anti-stereotypical—can mislead models into making a definitive choice when “UNKNOWN” is the correct answer. For example, the accuracy of Mistral-Inst. drops significantly from 66.91% in the w/o RAG setting to 45.49% with the stereo-set. This indicates a heightened risk of RAG systems producing seemingly confident but baseless answers in information-poor scenarios.
Conversely, for disambiguated questions, the impact of retrieval is more nuanced. While stereotypical documents often cause a minor decline in accuracy, the anti-set can be beneficial. Several models, including Llama3-8B-Inst. (60.52% to 66.07%) and Mistral-Inst. (67.26% to 71.88%), show improved accuracy with anti-stereotypical documents.
Regarding model scale, larger models generally exhibit higher accuracy. The accuracy of even the largest models, such as Qwen-72B-Inst., drops substantially on ambiguous questions when using the stereo-set (from 98.26% to 86.21%). The reasoning model, o4-mini, is notably more robust. On disambiguated questions, its accuracy remains exceptionally high and stable (96%) across all settings, indicating that its strong reasoning capabilities allow it to reliably prioritise the explicit evidence in the prompt over any distracting retrieved context.
These results indicate that, although retrieved documents might reduce accuracy in ambiguous questions, they can be beneficial in disambiguated settings when the retrieved documents offer relevant, counteracting signals against stereotypical biases.
§ F.3 Effect of the Retrievers on Accuracy
Table 14 compares the ambiguous and disambiguated accuracy scores for various models when retrieving documents from the full-set using three different retrieval methods: VectorIndex, BM25, and Contriever.
Among the retrieval methods, BM25 consistently yields higher ambiguous accuracy than both VectorIndex and Contriever. For instance, GPT-3.5 achieves an ambiguous accuracy of 39.93% with BM25, which is notably higher than the 27.58% obtained with VectorIndex and 31.80% with Contriever. Similar trends are evident for other models. In contrast, for the disambiguated questions the impact of the retrieval method is more varied. Some models such as Llama3-8B-Inst. and Qwen-14B-Inst., BM25 even lead to an improvement in the disambiguated accuracy relative to the w/o RAG setting.
§ F.4 Effect of Varying the Number of Retrieved Documents on Accuracy
Figure 7 and Figure 8 compare the accuracy of three LLMs under different numbers of retrieved documents. For the ambiguous questions, accuracy shows a general downward trend as more documents are retrieved. Because the retrieved texts are not directly relevant to the ambiguous query, and the additional information appears to introduce stereotypes (or anti-stereotypes) to the models, it can reduce the model’s ability to respond with “UNKNOWN”. By contrast, for the disambiguated questions, retrieving more documents sometimes achieve accuracy that is comparable (or at times exceeds) to the w/o RAG setting.
§ F.5 Multi-lingual Accuracy Evaluations
Table 15 and Table 16 present the accuracy for Chinese (CBBQ) and Japanese (JBBQ) datasets. In both of those languages, the highest ambiguous accuracy is achieved in the w/o RAG setting. When RAG is applied, the anti-set generally results in the lowest ambiguous accuracy. In contrast, for the disambiguated questions anti-set usually reports superior accuracy compared to the other RAG settings. These multilingual evaluations suggests a potential trade-off in RAG settings – ambiguous questions are best handled without retrieval or with a full-set corpus, whereas disambiguated questions benefit from retrieving documents from the anti-set, which also contributes to lower Diff-Bias scores.
Bias Category Setting GPT-3.5 Llama3-8B-Inst. Qwen2.5-7B-Inst. Qwen2.5-14B Qwen2.5-14B-Inst. Qwen2.5-72B-Inst.
Gender w/o RAG 45.24±2.17 / 75.74±1.87 50.03±2.18 / 60.52±2.13 81.45±1.70 / 52.03±2.18 63.79±2.10 / 82.84±1.65 96.53±0.80 / 71.92±1.96 98.26±0.51 / 68.25±2.03
stereo-set 27.83±1.96 / 71.68±1.97 26.19±1.92 / 63.74±2.10 62.60±2.11 / 53.72±2.18 46.83±2.18 / 77.63±1.82 87.05±1.47 / 68.30±2.03 86.21±1.38 / 68.61±2.02
full-set 27.58±1.95 / 73.12±1.94 24.36±1.87 / 63.89±2.10 62.95±2.11 / 56.10±2.17 49.75±2.18 / 77.08±1.83 89.78±1.32 / 67.76±2.04 89.44±1.23 / 65.97±2.07
anti-set 22.07±1.81 / 72.97±1.94 23.66±1.86 / 66.07±2.07 58.09±2.15 / 58.09±2.15 41.02±2.15 / 78.57±1.79 83.63±1.62 / 68.90±2.02 84.13±1.50 / 69.40±2.01
Age w/o RAG 18.97±1.27 / 88.70±1.02 31.03±1.49 / 75.57±1.39 60.08±1.58 / 77.42±1.35 42.80±1.60 / 92.53±0.85 78.76±1.32 / 89.24±1.00 84.41±1.48 / 91.93±1.14
stereo-set 16.63±1.20 / 81.55±1.25 16.03±1.19 / 70.03±1.48 48.64±1.61 / 77.99±1.34 35.30±1.54 / 89.65±0.98 75.95±1.38 / 83.32±1.20 73.18±1.78 / 91.50±1.20
full-set 19.97±1.29 / 83.21±1.21 15.76±1.18 / 71.47±1.46 51.17±1.62 / 79.59±1.30 40.84±1.59 / 90.43±0.95 87.93±1.05 / 84.51±1.17 84.62±1.48 / 90.08±1.29
anti-set 16.44±1.20 / 84.35±1.17 14.57±1.14 / 71.49±1.46 50.11±1.62 / 78.61±1.32 37.42±1.56 / 90.46±0.95 87.36±1.07 / 84.89±1.16 80.44±1.58 / 91.20±1.22
Race w/o RAG 56.97±2.24 / 83.62±1.67 59.04±2.22 / 77.55±1.89 94.04±1.07 / 68.03±2.11 80.85±1.78 / 93.62±1.10 98.94±0.46 / 78.46±1.86 99.10±0.27 / 83.83±1.58
stereo-set 33.88±2.14 / 83.24±1.69 35.00±2.16 / 78.03±1.87 70.32±2.07 / 75.85±1.93 58.35±2.23 / 92.66±1.18 91.33±1.27 / 83.83±1.66 93.78±0.81 / 82.55±1.63
full-set 35.74±2.17 / 85.16±1.61 37.61±2.19 / 78.14±1.87 73.40±2.00 / 76.06±1.93 62.71±2.19 / 94.04±1.07 94.04±1.05 / 84.15±1.65 96.23±0.69 / 80.00±1.74
anti-set 35.21±2.16 / 86.44±1.55 38.94±2.20 / 80.69±1.78 79.95±1.81 / 74.04±1.98 55.64±2.25 / 94.95±0.99 96.81±0.79 / 86.65±1.54 96.49±0.69 / 83.04±1.63
Religion w/o RAG 49.08±2.83 / 80.17±2.26 60.67±2.76 / 74.25±2.47 84.58±2.04 / 64.25±2.71 67.75±2.64 / 83.67±2.09 90.33±1.67 / 69.42±2.61 93.17±1.08 / 74.50±1.86
stereo-set 37.92±2.75 / 77.08±2.38 38.67±2.76 / 75.42±2.44 71.67±2.55 / 68.92±2.62 53.05±2.82 / 87.67±1.86 87.25±1.89 / 68.50±2.63 89.59±1.28 / 70.67±1.98
full-set 35.67±2.71 / 78.67±2.32 35.50±2.71 / 74.25±2.47 66.50±2.67 / 70.67±2.58 51.50±2.83 / 86.25±1.95 86.75±1.92 / 69.92±2.59 88.75±1.32 / 72.08±1.95
anti-set 30.50±2.61 / 78.58±2.32 32.92±2.66 / 76.25±2.41 67.17±2.66 / 71.75±2.55 47.67±2.83 / 87.92±1.84 84.42±2.05 / 72.17±2.54 87.34±1.37 / 76.09±1.86
Table 12: Accuracy scores for the ambiguous and disambiguated contexts (separated by ‘/’) for different bias categories and models, when document collections with varying degrees of social biases are used for retrieval. In each sub-category (Gender, Age, Race, Religion), the scores for each model are compared vertically with 95% CI half-widths shown as subscripts.
Model w/o RAG stereo-set full-set anti-set
GPT-3.5 45.24±2.17 / 75.74±1.87 27.83±1.96 / 71.68±1.97 27.58±1.95 / 73.12±1.94 22.07±1.81 / 72.97±1.94
Llama3-8B 25.94±1.91 / 41.96±2.15 21.03±1.78 / 47.97±2.18 21.43±1.79 / 47.82±2.18 20.78±1.77 / 49.40±2.18
Llama3-8B-Inst. 50.30±2.18 / 60.52±2.13 26.19±1.92 / 63.74±2.10 24.36±1.87 / 63.89±2.10 23.66±1.86 / 66.07±2.07
Mistral-Inst. 66.91±2.05 / 67.26±2.05 45.49±2.17 / 69.25±2.01 47.22±2.18 / 71.38±1.97 42.11±2.16 / 71.88±1.96
Llm-jp-1.8B 7.04±1.12 / 48.21±2.18 16.96±1.64 / 43.75±2.17 18.06±1.68 / 44.25±2.17 16.27±1.61 / 45.85±2.18
Llm-jp-3.7B 10.52±1.34 / 51.49±2.18 18.65±1.70 / 47.52±2.18 19.35±1.72 / 46.23±2.18 16.27±1.61 / 45.14±2.17
Llm-jp-13B 10.62±1.34 / 82.74±1.65 6.75±1.10 / 78.27±1.80 7.14±1.12 / 78.27±1.80 6.75±1.10 / 77.78±1.81
Qwen-3B 8.23±1.20 / 78.97±1.78 4.12±0.87 / 76.49±1.85 3.22±0.77 / 76.24±1.86 2.88±0.73 / 77.83±1.81
Qwen-3B-Inst. 68.20±2.03 / 58.43±2.15 57.19±2.16 / 57.14±2.16 52.28±2.18 / 63.10±2.11 54.37±2.17 / 59.42±2.14
Qwen-7B 30.65±2.01 / 67.31±2.05 26.49±1.93 / 68.40±2.03 26.74±1.93 / 69.54±2.01 24.31±1.87 / 69.05±2.02
Qwen-7B-Inst. 81.45±1.70 / 52.03±2.18 62.60±2.11 / 53.72±2.18 62.95±2.11 / 56.10±2.17 58.09±2.15 / 58.09±2.15
Qwen-14B 63.79±2.10 / 82.84±1.65 46.83±2.18 / 77.63±1.82 49.75±2.18 / 77.08±1.83 41.02±2.15 / 78.57±1.79
Qwen-14B-Inst. 96.53±0.80 / 71.92±1.96 87.05±1.47 / 68.30±2.03 89.78±1.32 / 67.76±2.04 83.63±1.62 / 68.90±2.02
Qwen-72B 70.49±1.73 / 83.29±1.59 62.21±1.71 / 83.98±1.56 63.84±1.82 / 82.29±1.64 57.05±1.88 / 85.12±1.51
Qwen-72B-Inst. 98.26±0.51 / 68.25±2.03 86.21±1.38 / 68.61±2.02 89.44±1.23 / 65.97±2.07 84.13±1.50 / 69.40±2.01
O4-mini 95.88±0.85 / 95.53±0.89 86.45±1.30 / 96.33±0.82 86.02±1.42 / 96.48±0.79 80.31±1.60 / 96.19±0.83
Table 13: Comparison of accuracy scores across different corpus settings on the BBQ gender dataset. Scores are reported in the format ambiguous / disambiguated, where higher values indicate better performance. For each model, the 95% CI half-widths are shown as subscripts
Model w/o RAG VectorIndex BM25 Contriever
GPT-3.5 45.24±2.17 / 75.74±1.87 27.58±1.95 / 73.12±1.94 39.93±2.14 / 74.21±1.91 31.80±2.03 / 73.07±1.94
Llama3-8B-Inst. 50.30±2.18 / 60.52±2.13 24.36±1.87 / 63.89±2.10 31.99±2.04 / 65.82±2.07 28.03±1.96 / 64.53±2.09
Qwen-7B-Inst. 81.45±1.70 / 52.03±2.18 62.95±2.11 / 56.10±2.17 60.32±2.14 / 57.29±2.16 61.61±2.12 / 54.86±2.17
Qwen-14B 63.79±2.10 / 82.84±1.65 49.75±2.18 / 77.08±1.83 45.88±2.18 / 84.52±1.58 47.57±2.18 / 78.27±1.80
Qwen-14B-Inst. 96.53±0.80 / 71.92±1.96 89.78±1.32 / 67.76±2.04 92.66±1.14 / 73.12±1.94 87.85±1.43 / 71.33±1.97
Table 14: Diff-Bias scores (ambiguous / disambiguated) using different retrieval methods from the full-set, with 95% CI half-widths as subscripts.
Figure 7: Accuracy for ambiguous questions for different numbers of retrieved documents.
Figure 8: Accuracy scores for disambiguated questions for different numbers of retrieved documents.
Bias Type Setting GPT-3.5-turbo Qwen-7B-Inst. Qwen-14B Qwen-14B-Inst. Qwen-72B-Inst.
Gender w/o RAG 26.52±2.81 / 64.30±3.05 90.48±1.87 / 45.88±3.17 72.62±2.84 / 57.30±3.15 96.32±1.20 / 40.84±3.13 93.29±0.96 / 56.90±2.14
stereo-set 13.31±2.16 / 67.08±2.99 37.55±3.08 / 64.09±3.05 42.42±3.15 / 60.29±3.11 76.73±2.69 / 45.78±3.17 71.21±1.66 / 55.35±2.16
full-set 16.77±2.38 / 67.28±2.99 43.40±3.16 / 62.04±3.09 49.46±3.18 / 61.11±3.10 84.52±2.30 / 48.97±3.18 85.39±1.35 / 54.22±2.17
anti-set 12.45±2.10 / 68.42±2.96 35.50±3.05 / 64.30±3.05 44.05±3.16 / 61.52±3.10 75.97±2.72 / 46.09±3.17 69.38±1.60 / 56.28±2.15
Age w/o RAG 7.08±1.15 / 65.66±2.04 83.69±1.55 / 64.55±2.07 49.61±1.75 / 75.98±1.86 87.21±1.39 / 66.41±2.05 77.20±1.70 / 85.79±1.51
stereo-set 13.53±1.35 / 57.37±2.15 43.46±2.13 / 72.66±1.95 41.70±1.83 / 73.93±1.92 76.91±1.79 / 50.83±2.18 66.06±1.95 / 76.52±1.85
full-set 11.24±1.28 / 60.69±2.12 41.60±2.11 / 75.20±1.89 39.89±1.79 / 77.44±1.82 82.03±1.62 / 57.28±2.15 73.54±1.79 / 79.01±1.77
anti-set 10.65±1.27 / 61.19±2.11 41.31±2.08 / 79.50±1.77 36.13±1.71 / 79.74±1.74 75.24±1.82 / 69.53±1.99 66.21±1.94 / 86.97±1.43
Race w/o RAG 61.24±1.93 / 58.70±2.11 98.97±0.34 / 37.37±2.10 73.77±1.30 / 65.30±2.05 99.87±0.16 / 49.23±2.17 99.81±0.14 / 50.06±2.18
stereo-set 50.78±2.17 / 56.51±2.11 64.65±1.96 / 56.06±2.15 55.60±1.54 / 60.42±2.10 94.08±0.70 / 41.47±2.14 92.19±0.99 / 39.91±2.12
full-set 48.63±2.18 / 61.20±2.08 63.42±1.96 / 58.46±2.13 59.51±1.68 / 63.28±2.07 94.40±0.69 / 45.90±2.16 93.69±0.72 / 45.51±2.16
anti-set 49.68±2.17 / 63.87±2.05 69.27±1.89 / 56.12±2.16 61.27±1.60 / 68.23±2.01 96.03±0.59 / 47.53±2.17 97.79±0.44 / 47.53±2.18
Religion w/o RAG 78.05±1.75 / 57.05±2.08 98.71±0.47 / 56.04±2.16 78.66±1.48 / 73.96±1.85 99.43±0.28 / 67.29±2.04 100.00±0.00 / 79.17±1.77
stereo-set 64.01±2.08 / 57.40±2.12 82.00±1.53 / 55.31±2.15 47.74±1.69 / 78.96±1.77 95.37±0.81 / 64.28±2.08 100.00±0.00 / 63.34±2.09
full-set 56.68±2.16 / 57.92±2.11 78.13±1.68 / 58.65±2.14 46.23±1.64 / 80.63±1.71 95.26±0.81 / 68.65±2.01 99.57±0.21 / 67.92±2.02
anti-set 58.94±2.14 / 60.83±2.09 79.96±1.68 / 54.90±2.14 41.81±1.76 / 84.38±1.57 94.40±0.87 / 67.50±2.03 97.42±0.47 / 70.84±1.98
Table 15: Accuracy scores for the CBBQ dataset (Chinese). Scores for ambiguous and disambiguated contexts are separated by ‘/’. For each model and bias type, the maximum accuracy is highlighted in bold.
Bias Type Setting GPT-3.5-turbo Qwen-7B-Inst. Qwen-14B Qwen-14B-Inst. Qwen-72B-Inst.
Gender w/o RAG 30.52±1.44 / 52.68±1.56 77.56±1.31 / 53.66±1.56 42.56±1.55 / 77.89±1.30 82.31±1.20 / 78.35±1.29 90.13±1.24 / 68.64±2.01
stereo-set 23.31±1.32 / 55.93±1.56 48.29±1.57 / 57.54±1.55 28.71±1.42 / 73.90±1.38 61.84±1.52 / 77.86±1.30 76.26±1.77 / 68.92±2.01
full-set 27.40±1.40 / 56.08±1.56 50.08±1.57 / 58.21±1.55 31.06±1.45 / 75.23±1.35 67.15±1.47 / 78.20±1.29 80.73±1.67 / 69.71±2.00
anti-set 24.26±1.34 / 56.29±1.55 45.35±1.56 / 58.00±1.55 25.95±1.37 / 77.28±1.31 61.61±1.52 / 80.52±1.24 75.72±1.78 / 71.14±1.97
Age w/o RAG 9.14±1.18 / 53.46±2.17 59.01±2.07 / 40.17±2.12 16.36±1.34 / 53.58±2.17 48.53±2.08 / 52.85±2.17 75.10±1.79 / 54.32±2.17
stereo-set 7.08±1.08 / 50.00±2.18 29.42±1.93 / 43.20±2.15 14.71±1.29 / 51.38±2.18 44.95±2.08 / 47.80±2.17 58.18±2.04 / 52.67±2.18
full-set 7.99±1.17 / 50.10±2.18 30.98±1.97 / 41.82±2.14 13.70±1.23 / 52.30±2.18 48.53±2.08 / 48.26±2.18 64.62±1.96 / 52.39±2.18
anti-set 7.00±1.06 / 52.57±2.18 33.18±2.01 / 43.29±2.15 12.78±1.16 / 51.75±2.18 46.33±2.06 / 49.91±2.18 66.82±1.94 / 53.68±2.17
Table 16: Accuracy scores for the JBBQ dataset (Japanese). Scores for ambiguous and disambiguated contexts are separated by ‘/’. For each model and bias type, the maximum accuracy is highlighted in bold.
Appendix G Detailed Data for Diff-Bias
This section show the detailed Diff-Bias scores for each separate model for the main part in the paper. Table 17, Table 18 and Figure 9 shows each model that contribute to the average score in Figure 3, Table 4 and Figure 4, respectively.
Bias Type Setting GPT-3.5 Llama3-8B-Inst. Qwen-7B-Inst. Qwen-14B Qwen-14B-Inst. Qwen-72B-Inst.
Gender w/o RAG 5.16±4.36 / -9.33±4.35 5.65±4.36 / 1.59±4.36 10.02±4.34 / -3.67±4.36 3.77±4.36 / -7.34±4.35 -2.38±4.36 / -2.38±4.36 -0.45±0.72 / 1.19±9.97
stereo-set 14.53*±4.32 / 7.14*±4.35 14.68*±4.32 / -0.40±4.37 24.01*±4.24 / 0.50±4.37 13.99*±4.32 / -2.68±4.36 4.61±4.36 / 2.68±4.36 12.21±2.01 / -0.50±10.07
full-set 11.31±4.34 / -0.10*±4.37 6.80±4.36 / -3.97±4.36 15.43±4.31 / -2.08±4.36 0.55±4.37 / -4.66±4.36 -3.08±4.36 / -5.95±4.36 5.80±1.83 / -4.76±9.87
anti-set 4.51±4.36 / -3.97±4.36 0.74±4.37 / -6.85±4.36 10.17±4.34 / -10.12±4.34 -4.51±4.36 / -8.93±4.35 -8.43±4.35 / -12.70*±4.33 0.79±2.35 / -5.56±10.12
Age w/o RAG 41.79±2.94 / 5.92±3.23 31.25±3.07 / 8.32±3.22 30.52±3.08 / 3.42±3.23 38.02±2.99 / 7.34±3.22 18.02±3.18 / 8.59±3.22 13.97±1.59 / 4.94±8.78
stereo-set 32.61±3.05 / 8.97±3.22 27.66±3.10 / 10.71±3.21 35.87±3.02 / 3.15±3.23 38.56±2.98 / 7.01±3.22 18.72±3.17 / 9.35±3.22 20.63±2.06 / 5.93±8.74
full-set 29.67±3.09 / 6.63±3.22 19.67±3.17 / 4.13±3.23 30.52±3.08 / 3.75±3.23 27.53±3.11 / 6.96±3.22 7.50±3.22 / 6.09±3.22 11.63±1.62 / 4.62±8.68
anti-set 17.83±3.18 / 6.30±3.22 8.97±3.22 / 2.77±3.23 20.11±3.16 / 3.53±3.23 6.96±3.22 / 6.79±3.22 2.69±3.23 / 3.26±3.23 8.64±1.87 / 4.79±8.73
Race w/o RAG 10.00±4.50 / 3.40±4.52 6.60±4.51 / 1.06±4.52 1.60±4.52 / 2.02±4.52 6.81±4.51 / 2.13±4.52 0.00±4.52 / -3.30±4.52 0.27±0.59 / 2.13±11.71
stereo-set 24.95*±4.38 / 13.30*±4.48 17.55*±4.45 / 9.26±4.50 12.55*±4.48 / 6.17±4.51 19.95*±4.43 / 8.09±4.51 3.88±4.52 / 3.83±4.52 2.50±1.26 / 2.77±11.62
full-set 16.60±4.46 / 8.83±4.50 12.18±4.49 / 6.91±4.51 7.98±4.51 / 4.89±4.51 13.46±4.48 / 3.83±4.52 0.00±4.52 / 0.64±4.52 -0.05±1.04 / 2.77±11.45
anti-set 6.49±4.51 / 5.43±4.51 7.23±4.51 / 4.15±4.52 4.73±4.52 / 6.11±4.51 4.36±4.52 / 2.23±4.52 -0.43±4.52 / -1.17±4.52 -1.38±1.03 / 0.53±11.66
Religion w/o RAG 8.92±5.64 / 4.33±5.65 18.76±5.56 / 7.17±5.64 5.92±5.65 / 3.50±5.65 12.58±5.61 / 5.00±5.65 8.17±5.64 / 2.83±5.66 5.83±2.02 / 4.33±13.81
stereo-set 14.83±5.60 / 12.50±5.61 17.67±5.57 / 8.50±5.64 16.67±5.58 / 5.83±5.65 22.67±5.51 / 7.17±5.64 10.42±5.63 / 9.33±5.63 8.25±2.43 / 5.00±13.46
full-set 8.00±5.64 / 9.00±5.64 10.17±5.63 / 9.17±5.63 12.83±5.61 / 5.50±5.65 12.58±5.61 / 8.83±5.64 8.42±5.64 / 4.17±5.65 5.75±2.58 / 4.17±13.59
anti-set 2.83±5.66 / 5.17±5.65 11.92±5.62 / 8.83±5.64 6.50±5.65 / 3.67±5.65 8.00±5.64 / 5.00±5.65 7.75±5.64 / 2.00±5.66 4.50±2.72 / 4.17±13.97
Table 17: Diff-Bias scores for the ambiguous and disambiguated contexts (separated by ‘/’) for different bias types and models, with document collections of varying social bias levels used for retrieval. In each bias type (Gender, Age, Race, Religion), the scores for each LLM are compared vertically (across the different settings). For each LLM and bias type, the maximum value of the ambiguous and disambiguated Diff-Bias scores are highlighted in bold red, while the minimum in bold blue (best viewed in colour). 95 % CIs that do not overlap with the corresponding w/o RAG setting are indicated by *.
Bias Type Setting GPT-3.5-turbo Qwen-7B-Inst. Qwen-14B Qwen-14B-Inst. Qwen-72B-Inst.
CBBQ
Gender w/o RAG 18.07±6.26 / 8.64±6.34 7.79±6.35 / 3.91±6.36 9.85±6.33 / 0.00±6.37 3.68±6.36 / 1.44±6.37 5.41±2.07 / 13.08±8.67
stereo-set 35.61±5.95 / 16.26±6.28 46.00±5.65 / 23.05±6.19 32.47±6.02 / 13.17±6.31 21.97±6.21 / 17.49±6.27 22.73±3.89 / 14.08±8.72
full-set 13.74±6.31 / 6.79±6.35 25.43±6.16 / 12.76±6.31 7.25±6.35 / 0.00±6.37 6.39±6.35 / -4.12±6.36 4.22±3.16 / 22.50±8.62
anti-set -6.39±6.35 / 6.38±6.35 -4.33±6.36 / -6.58±6.35 -6.60±6.35 / -10.70±6.33 -9.09±6.34 / -13.58±6.31 -15.26±4.25 / 18.42±8.62
Age w/o RAG 27.07±5.67 / 28.71±5.36 11.13±2.31 / -4.11±5.85 18.26±3.98 / -15.62±5.13 5.96±2.09 / -7.04±5.76 15.97±2.62 / -9.11±4.24
stereo-set 17.82±5.59 / 34.08±5.65 23.15±4.38 / 1.17±5.46 22.46±4.35 / -12.89±5.30 17.34±2.71 / -0.68±6.13 23.29±3.18 / 1.08±5.19
full-set 18.56±5.66 / 26.27±5.72 19.44±4.52 / 0.59±5.29 7.96±4.61 / -17.97±4.98 9.48±2.47 / -10.26±6.02 16.02±2.88 / -2.74±4.98
anti-set 11.04±5.75 / 32.33±5.58 1.56±4.68 / -9.18±4.92 1.96±4.78 / -15.53±4.80 -11.28±2.42 / -17.39±5.51 -4.69±3.45 / -7.72±4.04
Race w/o RAG 4.92±4.32 / 9.52±6.87 -0.13±0.59 / 18.96±6.99 4.23±2.86 / 7.44±6.61 -0.13±0.24 / -1.32±7.21 0.07±0.29 / -11.83±7.18
stereo-set 13.41±4.87 / -10.21±6.98 11.66±4.02 / -2.65±7.16 12.89±4.14 / 15.77±6.75 2.15±1.21 / 5.64±7.13 0.39±1.73 / -2.69±7.07
full-set 7.49±5.04 / -13.05±6.91 8.73±4.12 / -3.63±7.10 9.77±4.32 / 16.22±6.60 1.70±1.18 / 0.78±7.20 0.07±1.26 / -2.94±7.20
anti-set 12.70±4.94 / -10.37±6.84 3.00±3.82 / -4.66±7.19 8.14±4.92 / 11.87±6.43 1.37±1.00 / 2.18±7.21 0.39±0.75 / -2.18±7.22
Religion w/o RAG -2.80±4.14 / -5.63±8.54 1.29±1.02 / 24.17±8.60 6.46±3.71 / -16.67±7.48 0.43±0.42 / 17.92±8.19 0.00±0.00 / -11.67±7.19
stereo-set 5.61±5.44 / 4.38±8.67 14.34±3.38 / 5.21±8.83 24.68±5.83 / -13.75±7.14 4.64±1.69 / 1.88±8.51 0.00±0.00 / -13.34±8.54
full-set -3.88±5.99 / -2.08±8.62 9.16±4.01 / -2.29±8.78 13.04±6.22 / -21.25±6.78 2.16±1.75 / -1.46±8.28 0.43±0.42 / -9.17±8.30
anti-set -6.14±5.81 / -2.50±8.55 -3.88±3.95 / 6.88±8.78 5.39±6.71 / -13.75±6.37 -1.30±1.88 / 1.25±8.37 -2.59±1.01 / -8.34±8.09
JBBQ
Gender w/o RAG 1.51±3.13 / -4.75±3.13 1.53±3.13 / -5.06±3.13 8.77±3.12 / -16.00±3.09 -0.72±3.13 / -20.50±3.07 -3.58±1.34 / -2.82±4.11
stereo-set 12.17±3.11 / 2.15±3.13 13.11±3.11 / -7.21±3.13 17.41±3.09 / -9.36±3.12 11.84±3.11 / -19.22±3.08 1.56±2.10 / -2.97±4.10
full-set 11.50±3.11 / 1.84±3.13 10.35±3.12 / -5.98±3.13 11.58±3.11 / -12.93±3.11 4.22±3.13 / -15.59±3.10 -4.30±1.90 / -4.45±4.07
anti-set 3.25±3.13 / 0.31±3.13 10.53±3.12 / -4.65±3.13 9.56±3.12 / -17.94±3.08 3.73±3.13 / -19.79±3.07 -5.68±2.11 / -3.53±4.02
Age w/o RAG 24.57±7.82 / -6.18±8.37 17.47±5.15 / -1.74±8.25 35.48±7.07 / 3.26±8.38 28.68±5.64 / -2.95±8.38 17.01±3.81 / -1.57±8.37
stereo-set 15.63±7.99 / -3.68±8.40 22.61±6.79 / 1.76±8.33 29.04±7.36 / 4.51±8.39 24.36±5.85 / 3.83±8.39 27.30±4.82 / -4.78±8.39
full-set 13.97±7.97 / -5.35±8.39 16.91±6.83 / -0.38±8.30 23.44±7.55 / 1.14±8.40 14.16±5.86 / 1.41±8.41 18.29±4.63 / -4.94±8.39
anti-set 9.56±8.07 / -7.53±8.38 -0.09±6.86 / -1.58±8.34 4.32±7.84 / -1.04±8.40 5.15±6.08 / -2.39±8.41 10.76±4.64 / -3.56±8.38
Table 18: Diff-Bias scores for the CBBQ dataset (Chinese) and JBBQ dataset (Japanese). Scores for ambiguous and disambiguated contexts are separated by /’. For each model and bias type, the maximum value across settings is highlighted in bold red, and the minimum in bold blue.
Figure 9: Diff-Bias scores for ambiguous questions for different numbers of retrieved documents.
Figure 10: Diff-Bias scores for disambiguated questions for different numbers of retrieved documents.